# Visual Description Grounding Reduces Hallucinations and Boosts Reasoning in LVLMs

### v2

@@ Appendix I Model Details
! Model Methodology Helpfulness Clarity Factuality Depth Engagement LLaVA-v1 Vanilla-greedy 3.12 4.41 2.44 2.19 1.23 Vanilla-sampling 3.14 4.42 2.45 2.20 1.26 VDD 2.98 4.10 2.20 2.07 1.16 OPERA 3.20 4.48 2.50 2.22 1.28 Woodpecker 3.21 4.49 2.46 2.21 1.25 LRV 3.25 4.51 2.51 2.23 1.29 LURE 3.20 4.47 2.49 2.20 1.28 VDGD (ours) 3.42 4.59 2.78 2.30 1.30 LLaVA-1.5 Vanilla-greedy 3.25 4.41 2.66 2.61 1.19 Vanilla-sampling 3.24 4.40 2.65 2.58 1.10 VDD 2.82 4.10 2.44 2.39 1.20 OPERA 3.22 4.40 2.49 2.58 1.18 Woodpecker 3.26 4.42 2.46 2.64 1.22 LRV 3.25 4.38 2.52 2.59 1.24 LURE 3.24 4.35 2.51 2.60 1.21 VDGD (ours) 3.47 4.56 2.84 2.51 1.80 LLaVA-1.6 Vanilla-greedy 3.43 4.41 3.13 3.31 1.35 Vanilla-sampling 3.40 4.38 3.10 3.27 1.30 VDD 3.26 4.19 2.96 3.13 1.20 OPERA 3.49 4.48 3.20 3.38 1.40 Woodpecker 3.45 4.45 3.21 3.35 1.42 LRV 3.42 4.40 3.18 3.34 1.39 LURE 3.38 4.36 3.18 3.28 1.36 VDGD (ours) 3.64 4.71 3.49 3.42 1.45 mPLUG-Owl2 Vanilla-greedy 3.23 4.42 2.86 2.48 1.25 Vanilla-sampling 3.24 4.47 2.84 2.50 1.24 VDD 3.07 4.21 2.67 2.40 1.17 OPERA 3.29 4.52 2.92 2.58 1.27 Woodpecker 3.27 4.50 2.90 2.52 1.25 LRV 3.23 4.49 2.88 2.50 1.24 LURE 3.26 4.51 2.91 2.51 1.28 VDGD (ours) 3.34 4.64 3.15 2.70 1.32 InternLM-X Vanilla-greedy 3.28 4.52 3.35 2.59 1.37 Vanilla-sampling 3.27 4.53 3.33 2.57 1.35 VDD 3.08 4.24 3.10 2.44 1.30 OPERA 3.37 4.65 3.44 2.65 1.43 Woodpecker 3.38 4.63 3.45 2.66 1.44 LRV 3.40 4.69 3.47 2.70 1.46 LURE 3.36 4.64 3.42 2.68 1.42 VDGD (ours) 3.54 4.78 3.68 2.79 1.49 CogVLM Vanilla-greedy 3.58 4.51 3.35 2.76 1.38 Vanilla-sampling 3.59 4.50 3.37 2.78 1.41 VDD 3.36 4.32 3.21 2.58 1.30 OPERA 3.68 4.60 3.42 2.84 1.45 Woodpecker 3.70 4.58 3.45 2.85 1.42 LRV 3.65 4.57 3.40 2.82 1.40 LURE 3.72 4.62 3.46 2.83 1.41 VDGD (ours) 3.89 4.82 3.59 2.89 1.49 GPT-4-Turbo - 4.15 4.79 3.89 3.25 1.67

### v3

@@ Appendix I Model Details
! Model Methodology Helpfulness Clarity Factuality Depth Engagement LLaVA-v1 Vanilla-greedy 3.12 4.41 2.44 2.19 1.23 Vanilla-sampling 3.14 4.42 2.45 2.20 1.26 VCD 2.98 4.10 2.20 2.07 1.16 OPERA 3.20 4.48 2.50 2.22 1.28 Woodpecker 3.21 4.49 2.46 2.21 1.25 LRV 3.25 4.51 2.51 2.23 1.29 LURE 3.20 4.47 2.49 2.20 1.28 VDGD (ours) 3.42 4.59 2.78 2.30 1.30 LLaVA-1.5 Vanilla-greedy 3.25 4.41 2.66 2.61 1.19 Vanilla-sampling 3.24 4.40 2.65 2.58 1.10 VCD 2.82 4.10 2.44 2.39 1.20 OPERA 3.22 4.40 2.49 2.58 1.18 Woodpecker 3.26 4.42 2.46 2.64 1.22 LRV 3.25 4.38 2.52 2.59 1.24 LURE 3.24 4.35 2.51 2.60 1.21 VDGD (ours) 3.47 4.56 2.84 2.51 1.80 LLaVA-1.6 Vanilla-greedy 3.43 4.41 3.13 3.31 1.35 Vanilla-sampling 3.40 4.38 3.10 3.27 1.30 VCD 3.26 4.19 2.96 3.13 1.20 OPERA 3.49 4.48 3.20 3.38 1.40 Woodpecker 3.45 4.45 3.21 3.35 1.42 LRV 3.42 4.40 3.18 3.34 1.39 LURE 3.38 4.36 3.18 3.28 1.36 VDGD (ours) 3.64 4.71 3.49 3.42 1.45 mPLUG-Owl2 Vanilla-greedy 3.23 4.42 2.86 2.48 1.25 Vanilla-sampling 3.24 4.47 2.84 2.50 1.24 VCD 3.07 4.21 2.67 2.40 1.17 OPERA 3.29 4.52 2.92 2.58 1.27 Woodpecker 3.27 4.50 2.90 2.52 1.25 LRV 3.23 4.49 2.88 2.50 1.24 LURE 3.26 4.51 2.91 2.51 1.28 VDGD (ours) 3.34 4.64 3.15 2.70 1.32 InternLM-X Vanilla-greedy 3.28 4.52 3.35 2.59 1.37 Vanilla-sampling 3.27 4.53 3.33 2.57 1.35 VCD 3.08 4.24 3.10 2.44 1.30 OPERA 3.37 4.65 3.44 2.65 1.43 Woodpecker 3.38 4.63 3.45 2.66 1.44 LRV 3.40 4.69 3.47 2.70 1.46 LURE 3.36 4.64 3.42 2.68 1.42 VDGD (ours) 3.54 4.78 3.68 2.79 1.49 CogVLM Vanilla-greedy 3.58 4.51 3.35 2.76 1.38 Vanilla-sampling 3.59 4.50 3.37 2.78 1.41 VCD 3.36 4.32 3.21 2.58 1.30 OPERA 3.68 4.60 3.42 2.84 1.45 Woodpecker 3.70 4.58 3.45 2.85 1.42 LRV 3.65 4.57 3.40 2.82 1.40 LURE 3.72 4.62 3.46 2.83 1.41 VDGD (ours) 3.89 4.82 3.59 2.89 1.49 GPT-4-Turbo - 4.15 4.79 3.89 3.25 1.67
@@
Appendix M Computational Efficiency
Table 14 presents a computational analysis of various methods evaluated on LLaVA 1.5 using a 48GB GPU, comparing their average inference time and computational cost measured in teraFLOPs (T). While the VDGD method requires more computation than the baseline methods, the increase is modest and within a reasonable range when considering the significant performance gains it offers. Specifically, VDGD has an average inference time of 2.5 seconds and a computational cost of 11.9 teraFLOPs, which is a manageable increase compared to the Vanilla-greedy method’s 1.3 seconds and 9.3 teraFLOPs. Similarly, the combination of VDGD with a Small Captioning Model Ramos et al. (2023) demonstrates acceptable computational demands with an inference time of 2.1 seconds and 11.4 T. These results suggest that the enhanced performance of VDGD methods is achieved with computational requirements that are well within practical limits, making it a viable option.
Method Average Inference Time (s) FLOPs (T)
Vanilla-greedy 1.3 9.3
VCD 1.9 10.2
VDGD + Small Captioning Model 2.1 11.4
VDGD 2.4 11.9
Table 14: Comparison of average inference time and computational cost (in teraFLOPs) among VDGD and baseline methods evaluated on LLaVA 1.5 using a 48GB GPU.
Appendix N VDGD with Beam Search
VDGD can be seamlessly applied to beam search decoding without modifying its core methodology. Before a set of tokens is selected for each beam, VDGD reweights the logit space of the current logit based on the KL-Divergence between the current logit and the image description logits. This process is repeated at every decoding step. Another important point to note is that VDGD operates independently of prior response tokens, relying solely on the fixed image description tokens, which remain constant even during beam search. As a result, VDGD can be directly integrated into beam search decoding to achieve similar improvements in reducing hallucinations. Table 15 shows the result of VDGD with beam search on MMMU dataset.
Benchmark LLaVA-v1 LLaVA-1.5
Vanilla-greedy 1.26 1.35
Vanilla-sampling 1.27 1.44
VDD 1.34 1.52
OPERA 1.30 1.43
Woodpecker 1.32 1.44
LRV 1.29 1.49
LURE 1.31 1.47
PAI 1.42 1.39
HALC 1.40 1.54
VDGD 1.42 1.62
VDGD + Beam Search 1.39 1.58
Table 15: Performance comparison of VDGD w/ beam search for LLaVA-v1 and LLaVA-1.5
Appendix O Further Discussion on related works
Recent advances in understanding hallucinations in Large Vision-Language Models (LVLMs) provide critical insights into addressing the persistent challenge of hallucination in multimodal systems. Bai et al. (2024) highlight the growing importance of mitigating hallucinations in LVLMs, particularly in tasks involving complex interactions between visual and textual data. Their comprehensive survey dissects hallucinations into granular categories, such as object attributes and relations, and identifies factors contributing to hallucinations across data, model architecture, and inference stages.  Liu et al. (2023c) similarly focus on categorizing and analyzing hallucinations, emphasizing the importance of detailed evaluation metrics and benchmarks tailored to measure both generative and discriminative capabilities of LVLMs. They propose frameworks for systematically assessing hallucination symptoms and mitigating them through refined model alignment and enhanced multimodal representation. These studies underscore the necessity of bridging gaps in current LVLM evaluation methods and mitigation strategies by addressing unique challenges such as data quality, alignment module efficiency, and decoding mechanisms. Our proposed VDGD model aligns with these efforts by introducing a robust framework to reduce hallucinations, leveraging state-of-the-art insights while contributing novel mitigation pathways. This connection to the broader context enriches our understanding and opens pathways for integrating these findings into future iterations of VDGD and related LVLM frameworks.